OpenStudy (anonymous):

How do I factor x^4 + 1.

OpenStudy (anonymous):

Wolf gives me http://www.wolframalpha.com/input/?i=factor+x%5E4+%2B+1

OpenStudy (anonymous):

Which is what I need but I don't know how to get there.

OpenStudy (anonymous):

(x^2+1)(x^2-1) = (x^2 + 1) (x + 1)(x - 1)

OpenStudy (anonymous):

x^4+1 has no real roots, so at most you can factor it into complex roots using i^2=-1 x^4-i^2=(x^2)^2-(i)^2

OpenStudy (anonymous):

Someone suggested using the nth root thereom somehow, but that's all I know so far